a mothers age is three less than three times her daughters age. the sum of her age is 45. what is the mothers age?

Answer by algebrahouse.com(1659)

x = daughter's age
3x - 3 = mother's age {a mother's age is 3 less than 3 times her daughter's age}

x + 3x - 3 = 45 {sum of ages is 45}
4x - 3 = 45 {combined like terms}
4x = 48 {added 3 to each side}
x = 12 {divided each side by 4}
3x - 3 = 33 {substituted 12, in for x, into 3x - 3}

mother's age is 33

To ask a question, visit: www.algebrahouse.com